Let the modulus be p = 23. Bob sends to Alice the value B = 9. Alice’s secret value is a = 10. Compute the secret key kAB to be shared between Alice and Bob. (Explain your computation.) Tip: Perform computation of kAB as it is to be done by Alice. (For completeness, we note that the generator g = 6, but you do not need it for this computation.)
Answer #1

Solution:

=>Symmetric key cryptography is used.

=>Modulus(p) = 23

=>Bob's public key(B) = 9

=>Alice secret number(a) = 10

Explanation:

Calculating the secret key kAB:

=>Secret key(kAB) = (B)^a mod p

=>Secret key(kAB) = (9)^10 mod 23

(9^5 mod 23 = 8 so 8*8 mod 23 = 18)

=>Secret key(kAB) = 18

=>Hence the shared key between Alice and Bob = 18

I have explained each and every part with the help of statements attached to the answer above.